OpenStudy (lolipop_drum312):

How many days equal 5616 minutes? Enter your answer in the box. 5616 min = days (my answer 3.9 is it right)

OpenStudy (lolipop_drum312):

i'll be right back but still answer please

OpenStudy (mathmale):

You'll need to select and use several conversion factors. These stem from the following: 1 day = 24 hours 1 hour = 60 minutes

OpenStudy (rcp031):

1 day=24 hours=24*60 minutes =1440 minutes 1 day =1440 minutes that means 5616/1440=3.9 days or 3 days 9 hours precise

OpenStudy (lolipop_drum312):

so i was right

OpenStudy (mathmale):

Example: 1 mile= 5280 feet. Thus, the 2 possible conversion factors are 1 mile --------- 5280 feet or 5280 feet --------- 1 mile You wish to eliminate the unit of measurement "hours" in favor of the unit of measurement "days."

OpenStudy (lolipop_drum312):

so... um... it would be 1 mile?

OpenStudy (lolipop_drum312):

well thanks

OpenStudy (lolipop_drum312):

new Q and can you medal him too

OpenStudy (mathmale):

What I typed above are examples of two conversion factors.